OpenStudy (anonymous):

Can someone please help. Below are the steps to solve an equation: Step 1: |x − 4| + 1 = 8 Step 2: |x − 4| = 8 − 1 Step 3: |x − 4| = 7 Which of the following is a correct next step to solve the equation? x − 4 = −7 x + 4 = −7 −x − 4 = 7 x + 4 = 7

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you need to solve two equations to solve \[|x-4|=7\] one is \[x-4=7\] and the other is \[x-4=-7\]